alphaXtiger:

i need help 13*14+15-16= i am very stuck

Katrin369:

Okay if you're really stuck. What do we do first? I mean like in - * = /

0mega:

Use PEMDAS

supie:

\(\color{lime}{(13\times14)}+15-16=\) Fisrt solve 13*14=?

alphaXtiger:

ok

Katrin369:

@alphaxtiger wrote:
ok
Tell me what you got when you solve it.

alphaXtiger:

i got 182

Katrin369:

Tell me what you did. How did you solve it?

supie:

@katrin369 wrote:
@alphaxtiger wrote:
i got 182
Well, you're close.
they're right ...

supie:

Anyways now you're left with \(182+15-16\) do yk what you have to do next?

alphaXtiger:

yes

supie:

what do you have to do next ?

alphaXtiger:

i have to + 182+15???

supie:

yes, lmk what you get

alphaXtiger:

i got 197

supie:

Yes, so now you have 197-16=?

alphaXtiger:

it is 181

supie:

Yeah, that's your answer c:

alphaXtiger:

thx for your help

supie:

You're welcome.